You have mixed 100 g of 8% cream into 200 g of 3% cream. What is the resulting concentration?

Answer by checkley79(3341)

100*.08+200*.03=(100+200)X
8+6=300X
14=300X
X=14/300
X=.04667 OR 4.67% SOLUTION ANS.